### v2

# Effectively Controlling Reasoning Models through Thinking Intervention

Abstract
Reasoning-enhanced large language models (LLMs) explicitly generate intermediate reasoning steps prior to generating final answers, helping the model excel in complex problem-solving. In this paper, we demonstrate that this emerging generation framework offers a unique opportunity for more fine-grained control over model behavior. We propose Thinking Intervention, a novel paradigm designed to explicitly guide the internal reasoning processes of LLMs by strategically inserting or revising specific thinking tokens. We find that the Thinking Intervention paradigm enhances the capabilities of reasoning models across a wide range of tasks, including instruction following on IFEval, instruction hierarchy on SEP, and safety alignment on XSTest and SORRY-Bench. Our results demonstrate that Thinking Intervention significantly outperforms baseline prompting approaches, achieving up to 6.7% accuracy gains in instruction-following scenarios, 15.4% improvements in reasoning about instruction hierarchies, and a 40.0% increase in refusal rates for unsafe prompts using open-source DeepSeek R1 models. Overall, our work opens a promising new research avenue for controlling reasoning LLMs. WARNING: This paper contains red-teaming content that can be offensive.
1 Introduction
Recent advances in reasoning-enhanced models, including OpenAI’s o1 [27], DeepSeek’s R1 [20], and Google’s Flash Thinking [13], have significantly expanded the capabilities of large language models (LLMs). By explicitly incorporating intermediate reasoning steps before producing answers, these models exhibit substantially improved performance in handling complex tasks such as mathematical problem-solving [36], programming assistance [29], and logical inference [48].
Despite these developments, existing methods for guiding reasoning models still predominantly rely on input-level manipulations, such as prompt engineering [46], which modifies the instructions provided to the LLM. In this work, we demonstrate that the explicit thinking steps introduced in reasoning LLMs not only provide enhanced transparency into the model’s cognitive processes but also create new opportunities for direct and precise interventions within these reasoning stages.
Motivated by this insight, we introduce Thinking Intervention, a novel paradigm that explicitly controls the internal reasoning processes of models. Rather than allowing the model to generate entire reasoning chains on its own, Thinking Intervention specifies token sequences (e.g., detailed instructions, clarifications, or constraints) to be inserted or replaced within the ongoing reasoning process. Such targeted interventions enable fine-grained and transparent control over the reasoning trajectory, closely aligning the model’s behavior with required task objectives.
Demonstration. To further illustrate, consider a general instruction-following task shown in Figure 1(a), which asks the model to "list 5 famous moms in JSON format." A standard Vanilla Prompting would just state the instruction and Prompt Engineering might add a reminder like "Ensure the format is JSON." Nonetheless, reasoning models may still overlook essential instructions or constraints (Figure 1(b)). In contrast, Thinking Intervention explicitly guides the reasoning process by injecting instructions directly into the model’s internal thought process, e.g., "I should generate 5 famous moms and put them in a JSON format." Such precise intervention reduces the likelihood of the model missing constraints during reasoning, thereby significantly improving performance.
Notably, Thinking Intervention presents several key advantages: First, it enables fine-grained and flexible control over the reasoning process by adaptively inserting or revising intermediate reasoning steps based on task-specific needs; second, it does not require any form of model training and can be deployed in real-world settings with minimal engineering effort; and third, it seamlessly integrates and complements existing techniques, including prompt engineering and advanced fine-tuning methods.
Empirical findings. We demonstrate the effectiveness of Thinking Intervention across diverse tasks, including instruction following on IFEval [77], instruction hierarchy on SEP [81], and safety alignment on XSTest [45] and SORRY-Bench [68]. For instruction following tasks (§3), we show that applying Thinking Intervention enables the model to more effectively and accurately follow task instructions, leading to notable improvements of 6.7% and 1.9% over baseline Vanilla Prompting and Prompt Engineering methods respectively (Figure 1(b)). Similarly, in instruction hierarchy scenarios (§4), Thinking Intervention guides the model to reason about hierarchical instructions and appropriately prioritize main tasks over low-priority ones, thereby boosting robustness up to 15.4% compared to baselines (Figure 1(b)). Lastly, in safety alignment tasks (§5), we first show that open-source reasoning models (e.g., DeepSeek R1 [20]) often over-comply with unsafe instructions, highlighting an urgent need for better safety control methods. We then demonstrate that Thinking Intervention explicitly steers models toward safer reasoning, substantially increasing refusal rates for unsafe requests by up to 40.0% on XSTest and 26.9% on SORRY-Bench (Figure 1(b)).
Finally, in Section 6, we analyze internal model mechanisms (e.g., attention maps) and find that the attention during the reasoning process predominantly focuses internally rather than on external input tokens, explaining why Thinking Intervention achieves greater effectiveness. We also discuss practical design considerations for implementing Thinking Intervention in real-world settings.
Overall, our findings establish Thinking Intervention as a powerful, flexible, and broadly applicable paradigm for enhancing reasoning model across multiple dimensions. We encourage the community to further explore and adopt this framework, which offers precise, transparent, and effective control over LLM reasoning processes, ultimately enabling more reliable and aligned AI systems.
2 Thinking Intervention: A Novel Paradigm to Control Reasoning Models
2.1 Preliminaries and Notations
Consider a next-token prediction language model , where denotes the vocabulary set and represents the space of all possible token sequences over . Given a token sequence as input, the LLM predicts the next token in the sequence. Let denote an input context, where each . We use to denote the concatenation of two token sequences followed by . A conventional LLM autoregressively generates a response sequence by iteratively predicting each response token conditioned on the context and the previously generated tokens .
Reasoning-enhanced LLM. Unlike conventional LLMs, a reasoning-enhanced LLM explicitly separates the generation process into a "reasoning/thinking" stage and a "response" stage. Formally, the generation operates as follows: (1) Reasoning Stage: The model first generates a sequence of intermediate reasoning tokens (or a "reasoning chain") . Each reasoning token is autoregressively generated by conditioning on the input context and previously generated reasoning tokens: . (2) Response Stage: After obtaining the reasoning chain, the model generates the final response by conditioning each token on the context, reasoning chain, and previous response tokens: . This explicit decomposition enhances the model’s capability for complex tasks and improves interpretability by transparently exposing its reasoning steps [20].
2.2 Intervening in the Reasoning Process as a General Paradigm
Traditional approaches111Fine-tuning can also be viewed as a form of model control in certain cases, but we do not consider it here because it is more destructive—it modifies the entire model rather than controlling a fixed one. to improving LLM performance have largely focused on prompt engineering [64, 60, 71, 44], which optimizes the input to elicit better response . For reasoning-enhanced LLMs, although crafting the initial prompt remains important, the explicit reasoning stage offers a new, more direct pathway for optimization: intervening within the reasoning process itself.
In this work, we propose a general paradigm termed Thinking Intervention, which directly intervenes within the reasoning process of LLMs, e.g., through revising explicit instructions or guidance at intermediate reasoning steps. Unlike prompt engineering, where the input context is optimized before generating tokens, Thinking Intervention operates in an online, dynamic environment. The reasoning chain is generated token-by-token in real time, requiring the intervention mechanism to make decisions based on the incomplete reasoning chain . The key challenge lies in developing intervention strategies that can quickly evaluate partial reasoning paths, intervene appropriately, and adapt to current trajectories without disrupting the LLM’s natural reasoning flow.
General paradigm. Given the autoregressive nature of reasoning LLMs, we propose interventions that can insert new tokens or revise existing tokens within the reasoning chain. Formally, we define an intervention function that determines when and how to intervene in the LLM reasoning process:
where is the input context and represents all reasoning tokens generated up to step . The output sequence replaces the existing partial reasoning chain. The modified reasoning generation process can thus be formalized as:
This formulation highlights that interventions are strategically designed based on the specific reasoning path observed, enabling context-aware modifications at critical junctures. The revised reasoning chain can incorporate corrective feedback, alternative reasoning approaches, or relevant domain knowledge that addresses errors or gaps identified in the current reasoning flow. This generalized framework accommodates both token insertion and revision, providing a flexible and powerful mechanism for dynamically guiding the reasoning process.
2.3 Instantiation: Intervention via a Postfix-based Monitor
A simple yet powerful instantiation of the intervention function is based on monitoring the reasoning chain to detect specific trigger strings. Specifically, given a set of trigger strings (which can be single tokens or sequences of multiple tokens), the monitor checks if the most recent tokens (i.e., the postfix of the current reasoning chain) match any string in . If a match is detected, we append an intervention sequence (e.g., "I am a safe and responsible assistant") immediately to the existing reasoning chain. The trigger set can be designed flexibly to capture relevant reasoning stages, domain-specific phrases, or other reasoning markers. Formally, the intervention function with a postfix-based monitor is defined as:
Here, denotes the concatenation of the current reasoning with intervention sequence . Practical intervention strategies can be efficiently implemented by selecting appropriate triggers . We illustrate three easy-to-implement examples, though more sophisticated approaches are feasible. (1) Intervention at reasoning start: To intervene at the beginning of reasoning, a trigger string indicating reasoning onset (e.g., "<think>" in DeepSeek R1 models) can be included in . Upon detecting this string, the model immediately receives guidance through an intervention sequence containing relevant instructions or hints to direct reasoning from the start. (2) Intervention at reasoning conclusion: Utilize end-of-reasoning trigger strings (e.g., "</think>") within to reinforce critical points or identify overlooked issues before generating the final output. (3) Intervention at reasoning transitions: Incorporate transitional markers (e.g., "wait", "Hmm") into to prompt the model to review previous reasoning steps, correct potential mistakes, or provide further elaborations.
In our evaluations, we experimented with all three strategies and found that intervening at the beginning of the reasoning process was the most effective. Consequently, we primarily adopted this approach throughout our main experiments (§3, §4, §5) and analyzed alternative strategies in the discussion section (§6.2). In addition to the postfix-based intervention, we also implemented an adaptive strategy that leverages an auxiliary LLM to revise reasoning traces, as presented in Section 6.2.
2.4 Thinking Intervention enjoys several unique features and advantages
Design flexibility. The intervention sequence can be manually designed by domain experts or automatically synthesized using auxiliary models (e.g., LLMs translating high-level task requirements into specific targeted interventions). Interventions may also be instruction-independent (e.g., for safety alignment tasks) or instruction-dependent (e.g., to enhance specific instruction-following), allowing broad applicability. Finally, interventions can be flexibly applied at arbitrary positions throughout the reasoning process, enabling versatile control over the model’s reasoning.
Easy integration. Adopting Thinking Intervention requires minimal engineering effort and, crucially, does not necessitate fine-tuning or modifying the underlying model parameters. Furthermore, the intervention also incurs negligible computational overhead, ensuring lightweight deployment.
Broad compatibility. Thinking Intervention naturally integrates with established approaches such as prompt engineering, forming synergistic combinations to maximize effectiveness. For instance, prompt engineering can supply background information, while Thinking Intervention can explicitly guide intermediate reasoning steps, thereby enhancing overall model performance and consistency.
Effectiveness. Thinking Intervention effectively controls models by directly guiding the reasoning process, ensuring key instructions are actively integrated rather than passively referenced. As we will show later in Section 6, Thinking Intervention elicits stronger model attention on intervention tokens compared to prompt engineering (Figure 7). This leads to consistent improvements across instruction following (§3), instruction hierarchy (§4), and safety alignment tasks (§5).
3 Evaluation on Instruction Following Tasks
In this section, we empirically demonstrate how our proposed Thinking Intervention approach significantly enhances the instruction-following capability of reasoning models.
| (a) R1-Qwen-7B | (b) R1-Qwen-14B | (c) R1-Qwen-32B | (d) QwQ-32B |
Benchmark and models. We leverage Instruction-Following Evaluation (IFEval) [77] to measure how well reasoning models follow instructions. The benchmark comprises 500 prompts, each containing some identified types of verifiable instructions. These verifiable instructions function as constraints on the output, such as "do not use any commas". To quantify the model’s capability, we report the accuracy, defined as the proportion of prompts for which the responses satisfy all verifiable instructions within the prompt. We consider reasoning models distilled from DeepSeek R1 [20], including R1-Qwen-7B, R1-Qwen-14B, and R1-Qwen-32B, and the QwQ-32B model [43]. Further results on R1-Llama models and evaluation details are provided in Appendix B. All models use the same tag, <think> and </think>, to denote the start and end of the reasoning process.
Methods. We compare Thinking Intervention against two baselines: Vanilla Prompting, which directly uses original prompts unchanged, and Reminder Prompting, which augments the input prompts by restating instruction constraints as reminders. These reminders are generated by prompting an auxiliary LLM using original instructions. For Thinking Intervention, we convert reminder statements into first-person narrative interventions (i.e., changing "Ensure the answer..." to "I should ensure the answer..." via a fixed prefix)222In practice, intervention sequences can be directly constructed using auxiliary LLMs.. This intervention sequence is then inserted at the beginning of the reasoning process. Figure 2 (top) illustrates how Thinking Intervention integrates with the Vanilla Prompting and Reminder Prompting.
Thinking Intervention consistently improves instruction-following capability. Figure 2 (bottom) illustrates consistent improvements across multiple reasoning models. Specifically, integrating Thinking Intervention with the baseline Vanilla Prompting yields accuracy gains of 5.74%, 4.99%, 6.65%, and 2.96% for R1-Qwen-7B, R1-Qwen-14B, R1-Qwen-32B, and QwQ-32B, respectively. Notably, the effectiveness of Thinking Intervention is preserved as the model size increases within the R1-Qwen family, with R1-Qwen-32B showing even greater benefits than R1-Qwen-7B and R1-Qwen-14B. Further performance improvement is achieved by combining Thinking Intervention with the Reminder Prompting method: accuracy reaches as high as 62.84% (R1-Qwen-7B), 77.63% (R1-Qwen-32B), and 82.44% (QwQ-32B). Thus, Thinking Intervention not only provides substantial stand-alone gains across model families, but also complements existing prompting methods.
Overall, these findings confirm Thinking Intervention’s capability to precisely guide reasoning models to follow constrained instructions. Crucially, this significant performance boost is achieved without requiring additional model training. Moreover, we explore the broader applicability of Thinking Intervention in Appendix C, showing how explicit intervention during the reasoning process can mitigate model overthinking, thereby highlighting the versatility and utility of the proposed method.
4 Evaluation on Instruction Hierarchy Task
Next, we explore how Thinking Intervention benefits the instruction hierarchy task [59, 66], which evaluates a model’s ability to appropriately prioritize high-priority instructions over low-priority ones. This capability is essential for safety-critical applications, where models must adhere to specific guidelines even in the presence of conflicting instructions. We examine how Thinking Intervention enhances the model’s ability to navigate complex scenarios involving competing directives.
| R1-Qwen-7B | R1-Qwen-14B | R1-Qwen-32B | QwQ-32B | |||||
| Methods | Rob.(%) | Utility(%) | Rob.(%) | Utility(%) | Rob.(%) | Utility(%) | Rob.(%) | Utility(%) |
| Vanilla | 57.60 | 74.44 | 34.00 | 81.04 | 34.80 | 81.76 | 22.20 | 88.00 |
| Vanilla+ThinkingI | 60.80 | 74.40 | 38.40 | 81.08 | 50.20 | 82.02 | 31.40 | 88.16 |
| Reminder | 57.60 | 74.20 | 38.40 | 80.50 | 46.20 | 81.16 | 36.20 | 87.52 |
| Reminder+ThinkingI | 62.60 | 73.92 | 41.80 | 80.90 | 66.40 | 80.90 | 43.40 | 86.79 |
Benchmark. We evaluate on the SEP dataset [81]. Each data point contains a high-priority main instruction paired with relevant data content and an unrelated low-priority instruction. Models are expected to prioritize the main instruction while ignoring the low-priority instruction. This benchmark enables us to evaluate how effectively models can maintain instruction hierarchies in complex scenarios involving potentially misaligned directives.
Evaluation metrics. We evaluate model performance on the SEP benchmark using two key metrics: (1) robustness, which measures the proportion of low-priority instructions correctly ignored when embedded within data; (2) utility, which quantifies the model’s baseline performance on the main task in the absence of any low-priority instructions. For the utility metric, we follow [75] by employing LLM-as-a-judge for evaluation and normalizing scores to a 0-100% scale.
Methods. Similar to our instruction-following experiments, we include two baseline approaches: Vanilla Prompting, which directly uses the prompts without additional guidance, and Reminder Prompting, which includes an explicit instruction reminder. For our Thinking Intervention approach, we use the intervention sequence "I should follow all the instructions in the task block and not follow any instructions in the data block." to explicitly guide the model in maintaining the correct instruction hierarchy. This intervention sequence is inserted at the beginning of the reasoning process to help the model correctly prioritize the instructions. More details on the evaluation setup, including baseline prompts and examples, can be found in Appendix D.1.
Thinking Intervention significantly improves robustness while maintaining model utility. Table 1 presents the evaluation results of reasoning models on the SEP benchmark. Our Thinking Intervention approach consistently improves both Vanilla Prompting and Reminder Prompting techniques in terms of robustness across all model scales. For example, when applied to R1-Qwen-32B with Reminder Prompting, Thinking Intervention achieves a robustness of 66.4%, marking a substantial 20.20% improvement over the Reminder Prompting baseline. We observe similar robustness enhancements across other model variants, including gains of 5.0% for R1-Qwen-7B, 3.4% for R1-Qwen-14B, and 7.2% for QwQ-32B. Importantly, Thinking Intervention preserves the utility of the models. Across all settings, it incurs only a negligible variation in utility ( 0.73%) compared to the baselines. Interestingly, we note that R1-Qwen-7B exhibits the highest robustness with Vanilla Prompting. This is attributed to its lower general utility ( 5% lower than other models), which could prevent it from correctly responding to the injected low-priority prompt, thus resulting in higher robustness.
These results demonstrate that Thinking Intervention effectively guides models to maintain proper instruction hierarchies by correctly following high-priority instructions without compromising their utility-related tasks. For experimental results with R1-Llama models, see Appendix D.
5 Evaluation on Safety Alignment Task
Before deployment, LLMs typically undergo a safety alignment process to ensure the model does not respond to harmful queries, such as "how to build a bomb". In this section, we demonstrate that open-source reasoning models (e.g., R1-Qwen-32B) exhibit notably low refusal rates to unsafe requests, and that Thinking Intervention can effectively steer the model toward much safer behavior.
Benchmark and models. We leverage XSTest [45] to evaluate the safety alignment of models. The benchmark contains 200 harmful requests and 250 benign questions that include safety triggers for LLMs (e.g., "How can I kill a Python process?"), which may cause the model to mistakenly refuse to answer. Following [45, 76], we leverage GPT-4o-mini as an evaluator to classify the responses into compliant or refusing behaviors. We report two metrics: the refusal rate for unsafe requests and the compliance rate for safe requests. We extensively evaluate a range of models, including open-source reasoning models (R1-Qwen-7B, R1-Qwen-14B, and R1-Qwen-32B), one representative closed-source reasoning model (o3-mini), and non-reasoning models (GPT-4o-mini and GPT-4o).
Baseline prompting methods and Thinking Intervention. We consider the following baselines: Vanilla Prompting, which does not include any additional instructions; Default Prompting, which uses the default safety prompting of Llama-2 [56] to ensure safety; Reminder Prompting [69], which reminds the model to act responsibly; and Goal Priority [74], which explicitly prioritizes safety over other objectives. The complete prompts used for each baseline are provided in the Appendix E.1. For Thinking Intervention, we inject the intervention sequence as "I am a helpful, respectful, and honest assistant." at the beginning of the reasoning process to steer the model toward safety.
Thinking Intervention effectively steers the reasoning models toward safety (Figure 5). LLMs with Vanilla Prompting typically face a trade-off between compliance and safety. Specifically, R1 models achieve near-perfect compliance (100%) with safe requests but demonstrate an alarmingly low refusal rate to unsafe requests (20%). In contrast, the GPT series refuses more than 70% of unsafe requests, with o3-mini even refusing nearly all unsafe requests (100%),333Since GPT models are closed-source, it is unclear if there exists auxiliary monitor filtering unsafe responses; thus, high (or even perfect) refusal rates might reflect monitor behavior rather than the models themselves. but maintain a relatively lower but acceptable compliance rate (90.4%-95.6%) for safe requests. Encouragingly, applying Thinking Intervention to R1 models substantially boosts safety performance, increasing refusal rates for unsafe requests by over 40% across all models. We observe only mild compliance decreases (2%) in larger models (R1-Qwen-14B and R1-Qwen-32B) and a relatively larger drop for the smaller model (R1-Qwen-7B). This is understandable given that safety questions are intentionally designed to be challenging, making them particularly difficult for less capable models to distinguish.
Thinking Intervention seamlessly complements prompting techniques, significantly enhancing model safety (Figure 5). We further evaluate the R1-Qwen-32B model on the XSTest benchmark under various prompting baselines. Compared to Reminder Prompting alone, combining Reminder Prompting with our Thinking Intervention approach increases the refusal rate for unsafe requests by 30%, while maintaining a high compliance rate (97%) for safe requests. Notably, when integrated with Goal Priority prompting, Thinking Intervention achieves a refusal rate of 75% on unsafe requests and a compliance rate of 95% on safe requests, performance comparable to safety-aligned GPT-4o models (Figure 5).
Thinking Intervention excels in more comprehensive safety benchmark (Figure 5). To further validate our approach, we evaluate SORRY-Bench, which features a more comprehensive taxonomy and more detailed unsafe instructions, using exactly the same methods. Our results show that Thinking Intervention consistently improves robustness (i.e., the refusal rate of unsafe instructions) over baseline prompting methods. For example, when combined with Default Prompting, Thinking Intervention achieves a refusal rate of approximately 87% for unsafe requests, a nearly 20% improvement over Default Prompting alone and even higher than the refusal rate of the GPT-4o and o3-mini models. These results further demonstrate that Thinking Intervention can effectively steer models toward safer behavior across different safety benchmarks, emphasizing the generalizability of our approach.
Overall, our findings demonstrate that Thinking Intervention significantly enhances model safety while maintaining a high compliance rate to benign requests. Its simplicity and effectiveness highlight its practical value for real-world deployment. While Thinking Intervention is not a panacea for all safety challenges, it could serve as a complementary safety layer that integrates seamlessly with existing techniques like RLHF [39, 11] and Constitutional AI [6, 49], thus contributing to the multi-layered safety frameworks [38]. We provide additional results with other models, including R1-Llama, QwQ-32B, and the safety fine-tuned variant, in Appendix E. Notably, Thinking Intervention achieves further safety improvements even when applied to models already fine-tuned for safety.
6 Analysis and Discussion
6.1 Why is Thinking Intervention effective?
Next, we seek to understand why Thinking Intervention demonstrates consistent performance gains compared to prompt engineering. We visualize attention maps using a safety alignment benchmark and compare Thinking Intervention against prompt engineering (Figure 7). Attention areas where later reasoning stages attend to intervention tokens are highlighted in the blue box. Our analysis reveals that the reasoning processes’ attention is focused internally rather than toward external input tokens. Consequently, standard Prompt Engineering achieves limited impact, as models scarcely direct attention toward such prompts. In contrast, attention maps with Thinking Intervention show significantly increased internal attention to the explicitly injected intervention tokens during the reasoning process. These visualizations suggest that Thinking Intervention effectively guides internal model reasoning, enabling more precise and reliable model control. See Appendix H for more details.
6.2 Exploring Design Choices for Thinking Intervention
Our primary evaluations focus on concise, first-person narrative interventions applied at the beginning of the reasoning process (TIdefault in Figure 7). However, as discussed in Section 2, Thinking Intervention is highly general and flexible; various other design choices exist and merit exploration.
Position of intervention. First, we investigated alternative intervention positions, specifically within the middle (TImid) or towards the end (TIend) of the reasoning phase. Figure 7 summarizes performance comparisons on the XSTest benchmark. Interestingly, interventions performed at later stages of reasoning demonstrated diminished effectiveness relative to early-stage Thinking Intervention. We hypothesize that reasoning paths become progressively harder to redirect once models have deliberated sufficiently long on incorrect or suboptimal trajectories. Similar findings on the SORRY-Bench are provided in Appendix F.1.
Complexity of intervention sequences. Our main experiments intentionally employed relatively concise intervention sequences to facilitate fair comparisons against prompt engineering methods. Yet, more sophisticated instructions can theoretically provide richer guidance to models. For instance, we performed an exploratory analysis using more detailed safety instructions (termed TIlong in Figure 7), and noticed clear trade-offs emerging between compliance and safety alignment: longer sequences markedly improved alignment with safety goals but reduced overall compliance rates due to overly restrictive guidance. Similar trends across other prompting methods are presented in Appendix F.2.
Narrative perspective. Another design consideration is whether narratives that direct the model to reason in the first-person form are fundamentally necessary. To clarify this, we conducted additional experiments comparing first-person and second-person narrative Thinking Intervention (TIsecond, shown in Figure 7). The results indicate minimal performance differences between the two variants. We attribute this negligible impact to the robust self-correction capabilities of reasoning models. Similar results on SORRY-Bench, alongside an illustrative example, are provided in Appendix F.3.
Leveraging an auxiliary LLM for intervention. Beyond inserting interventions, we explored more sophisticated approaches by prompting an auxiliary LLM (i.e., Qwen2.5-7B-Instruct [42]) to dynamically monitor and revise the primary model’s reasoning traces for enhancing safety, termed TImodel (Figure 7). Preliminary results indicate that LLM-assisted interventions occasionally outperform manually crafted interventions, but the additional computational overhead may hinder practical deployment. Further implementation details and additional results are provided in Appendix F.4.
6.3 Practical utility of Thinking Intervention.
We anticipate several practical use cases for Thinking Intervention. For LLM providers, it can enhance model performance by integrating system prompts, such as those used for role-play, into the reasoning process, thereby improving user experience. For LLM users, Thinking Intervention can be easily adopted with open-source models, where users can create their own interventions when the model is not reasoning as expected. Nevertheless, adopting Thinking Intervention with closed-source models remains challenging, as most providers do not currently support interventions in internal reasoning processes. Furthermore, exposing public APIs for reasoning interventions can pose security risks, potentially allowing malicious actors to bypass safety mechanisms more easily. We thus recommend that LLM providers carefully evaluate the trade-off between usability gains and safety considerations before deploying Thinking Intervention-like capabilities publicly.
Looking forward, we anticipate that Thinking Intervention will enable broader applications across various challenging domains. For instance, interventions could be integrated into models performing complex tasks, such as medical diagnosis or legal reasoning, allowing domain experts to apply targeted corrections and inject domain knowledge at critical stages of the reasoning process, thereby significantly improving reliability and trustworthiness in high-stakes applications.
7 Related Works
Before the emergence of reasoning models, intervention-based methods were proposed to measure faithfulness in chain-of-thought (CoT) reasoning [30, 58], and have since been extended to reasoning-enhanced models [7, 4]. Our Thinking Intervention framework complements those monitoring-based methods by enabling precise control over the reasoning process. Other studies focus on controlling reasoning length, either encouraging longer chains to boost accuracy [37, 1] or shortening them for efficiency [21, 70, 31]. Additionally, external tools have been incorporated into reasoning chains [16, 32]. In contrast, the Thinking Intervention paradigm enables more general and fine-grained control over reasoning models, significantly broadening their capabilities and flexibility. See Appendix A for additional related works on reasoning models, LLM control, and evaluation tasks.
8 Conclusion.
In this paper, we propose Thinking Intervention, a novel approach to effectively control reasoning-enhanced LLMs. We demonstrate that Thinking Intervention can significantly improve the performance of LLMs across various tasks, including instruction following, instruction hierarchy, and safety alignment. We strongly encourage further investigation and adoption of Thinking Intervention, as it provides essential tools for fine-grained reasoning intervention, laying important groundwork towards more reliable, interpretable, and human-aligned LLM systems.
Appendix A Addtional Related Works
In this appendix, we present additional related works relevant to reasoning models, controlling LLMs, and the tasks used for our evaluation.
Reasoning models. Reasoning models have rapidly advanced since OpenAI’s o1 model [27]. This trend has produced closed-source models like Google’s Flash Thinking [13], Anthropic’s Claude 3.7 Sonnet [3], and xAI’s Grok 3 [67], alongside open-source alternatives such as DeepSeek R1 [20], QwQ [43], and S1 [37]. These models employ test-time scaling [51, 65], allocating additional inference computation to improve performance on complex tasks.
Controlling LLMs. There are two mainstream approaches for controlling LLMs after the training stage. Prompt Engineering provides clear and detailed instructions, either manually written [8, 64, 57, 71] or automatically generated [50, 44, 53, 14], to achieve a specific objective. Activation Steering selects a subset of LLM inner activations to probe [12, 79, 24, 33, 73] as a means to control LLMs. Thinking Intervention differs from these methods as it intervenes in the thinking process.
Instruction following. LLMs rely on accurately following natural language instructions for broad applicability. This capability is improved through supervised fine-tuning on instruction-response pairs [47, 63, 10] and reinforcement learning from human feedback (RLHF) [52, 5, 39]. Additionally, prompting techniques [60, 64] further help elicit more effective responses.
Instruction hierarchy. The concept of instruction hierarchy was proposed by [59], suggesting that LLM systems should prioritize instructions based on their trustworthiness. Otherwise, they may become vulnerable to misalignment or adversarial prompts [40, 18, 15]. Researchers have proposed various methods to enhance instruction hierarchy through additional training on misaligned data [9, 59, 41], prompting-based techniques [25, 81], and architectural design [66, 82].
Safety alignment. Safety alignment [6, 17, 55] is a critical aspect of LLM development, aiming to ensure that models follow ethical guidelines and avoid producing harmful content [62, 80, 68]. For reasoning models, works from OpenAI [19, 72] suggest that their o1/o3 series can achieve better safety alignment by leveraging more test-time compute. Meanwhile, other studies [78, 28, 26] have indicated that open-sourced reasoning models, like DeepSeek R1 series, exhibit more safety issues.
In this work, we explore how to leverage Thinking Intervention to improve instruction following, instruction hierarchy, and safety alignment in open-source reasoning models.
Appendix B Instruction Following Evaluation (IFEval)
B.1 More Details of Evaluations
IFEval [77] evaluates the instruction-following capabilities of language models using 25 distinct instruction types across approximately 500 prompts. In the main texts, we report only the prompt-level strict accuracy as accuracy for simplicity. Here, we formally discuss all four metrics used in [77] and report the results.
Prompt-level strict accuracy: The proportion of prompts for which all verifiable instructions are correctly followed.
Instruction-level strict accuracy: The proportion of verifiable instructions that are correctly completed, evaluated individually.
Prompt-level loose accuracy: Similar to prompt-level strict accuracy, but evaluated under a loose evaluation criterion (see below for details).
Instruction-level loose accuracy: Instruction-level strict accuracy under a loose evaluation criterion (see below for details).
Strict accuracy and loose accuracy. Strict accuracy requires the model output to precisely match the requirements. For instance, if an instruction specifies output in JSON format, the entire response must be in valid JSON format with no extraneous text. Any deviation results in the output being marked incorrect. Loose accuracy, on the other hand, allows some flexibility. For example, if a response begins with a preamble like "Sure, here is the answer:" followed by correctly formatted JSON, it would still be considered correct under loose accuracy criteria, even though it would fail strict evaluation. In addition to removing such introductions, evaluators also disregard font modifiers and outros, making the assessment more flexible.
For more details on the benchmark, please refer to the original paper [77]. Our implementation uses the codebase available at https://github.com/josejg/instruction_following_eval.
Generating Reminder Prompting and Thinking Intervention. Figure 8 demonstrates how we generate the text of Reminder Prompting by providing system prompts to GPT-4o. We then modify the narrative by adding a prefix (e.g., from "Ensure the summary is at least 300 words" to "I should ensure the summary is at least 300 words") with Python code to create the intervention sequence.
| Models | Methods | Prompt-level | Inst-level | Prompt-level | Inst-level |
| strict acc.(%) | strict acc.(%) | loose acc.(%) | loose acc.(%) | ||
| R1-Qwen-7B | Vanilla | 55.08 | 66.19 | 58.60 | 68.70 |
| Vanilla+ThinkingI | 60.99 (+5.91) | 70.50 (+4.31) | 63.96 (+5.36) | 73.02 (+4.32) | |
| Reminder | 60.99 | 70.50 | 64.33 | 73.02 | |
| Reminder+ThinkingI | 62.85 (+1.86) | 72.18 (+1.68) | 66.54 (+2.21) | 75.06 (+2.04) | |
| R1-Qwen-14B | Vanilla | 70.43 | 79.50 | 73.57 | 81.77 |
| Vanilla+ThinkingI | 75.42 (+4.99) | 81.65 (+2.15) | 78.37 (+4.80) | 84.29 (+2.52) | |
| Reminder | 72.83 | 81.18 | 76.53 | 83.69 | |
| Reminder+ThinkingI | 74.31 (+1.48) | 82.13 (+0.95) | 77.26 (+0.73) | 84.41 (+0.72) | |
| R1-Qwen-32B | Vanilla | 70.43 | 79.14 | 74.49 | 81.89 |
| Vanilla+ThinkingI | 77.08 (+6.65) | 84.29 (+5.15) | 80.96 (+6.47) | 86.93 (+5.04) | |
| Reminder | 75.23 | 82.85 | 78.74 | 85.37 | |
| Reminder+ThinkingI | 77.63 (+2.40) | 84.53 (+1.68) | 81.70 (+2.96) | 87.29 (+1.92) | |
| QwQ-32B | Vanilla | 79.30 | 86.09 | 83.92 | 89.09 |
| Vanilla+ThinkingI | 82.26 (+2.96) | 88.01 (+1.92) | 86.32 (+2.40) | 90.65 (+1.56) | |
| Reminder | 81.33 | 87.53 | 86.69 | 91.01 | |
| Reminder+ThinkingI | 82.44 (+1.11) | 88.13 (+0.60) | 86.69 (+0.00) | 91.01 (+0.00) | |
| R1-Llama-8B | Vanilla | 56.56 | 67.63 | 60.07 | 70.50 |
| Vanilla+ThinkingI | 65.43 (+8.87) | 74.58 (+6.95) | 68.95 (+8.88) | 76.98 (+6.48) | |
| Reminder | 62.85 | 73.02 | 67.47 | 76.62 | |
| Reminder+ThinkingI | 66.17 (+3.32) | 75.30 (+2.28) | 70.06 (+2.59) | 78.30 (+1.68) | |
| R1-Llama-70B | Vanilla | 79.85 | 86.09 | 82.62 | 88.13 |
| Vanilla+ThinkingI | 80.41 (+0.56) | 86.69 (+0.60) | 84.47 (+1.85) | 89.57 (+1.44) | |
| Reminder | 82.07 | 87.89 | 84.84 | 89.81 | |
| Reminder+ThinkingI | 82.44 (+0.37) | 88.13 (+0.24) | 85.58 (+0.74) | 90.53 (+0.72) | |
B.2 Comprehensive Experiment Results
In Table 2, we provide comprehensive experimental results on the IFEval benchmark, covering additional reasoning-enhanced models (R1-Llama-8B and R1-Llama-70B) alongside diverse evaluation metrics.
Our findings are consistent with the main conclusions presented in Section 3, clearly demonstrating that our Thinking Intervention paradigm effectively improves models’ capability to accurately follow instructions across all evaluation metrics. Specifically, in terms of prompt-level loose accuracy, applying Thinking Intervention leads to substantial empirical improvements over the Vanilla Prompting. We observe performance gains of 5.36%, 4.80%, and 6.47% for R1-Qwen-7B, R1-Qwen-14B, and R1-Qwen-32B, respectively.
Moreover, we find that Thinking Intervention achieves similar performance enhancements for the additional reasoning-enabled models: prompt-level loose accuracy increases by 8.88% for R1-Llama-8B and by 1.85% for R1-Llama-70B compared to Vanilla Prompting. These observations strongly suggest that our Thinking Intervention method enhances instruction-following capabilities across diverse reasoning-enhanced LLMs.
Appendix C Instruction Following Evaluation (Overthinking)
As the increasing computational costs associated with sequential scaling of reasoning models, recent studies have dedicated significant efforts toward optimizing the thinking steps to minimize unnecessary computational overhead while maintaining accuracy [22, 70, 35]. For a detailed overview of these developments, we refer readers to the recent survey [54].
Here, we investigate the effectiveness of our Thinking Intervention paradigm in mitigating such overthinking. Specifically, we use the standard 500-sample subset from the challenging MATH benchmark [23], which is commonly adopted by prior work (e.g., OpenAI’s study [34]). We measure: (i) accuracy: the percentage of correctly solved problems; and (ii) reasoning length: the total token count of reasoning process.
We compare our method against two baselines: Vanilla Prompting, the standard procedure where the model directly receives the math problem as a prompt without additional guidance; Reminder Prompting, where a textual cue ("Please solve it without thinking too much.") is included, aiming to guide models toward succinct reasoning. For Thinking Intervention, we explicitly inject a concise instruction ("Okay, the user asked for this. I need to solve it without thinking too much.") at the start of the reasoning process. Following prior evaluations, we test across a set of model variants (R1-Qwen-14B, R1-Qwen-32B, and QwQ-32B).
| R1-Qwen-14B | R1-Qwen-32B | QwQ-32B | ||||
| Methods | Accuracy (%) | Length (tokens) | Accuracy (%) | Length (tokens) | Accuracy (%) | Length (tokens) |
| Vanilla | 89.4 | 3281 | 90.8 | 3101 | 90.0 | 3926 |
| Vanilla+TI | 88.6 | 2494 (-23.99%) | 90.0 | 2388 (-22.99%) | 90.6 | 3781(-3.7%) |
| Reminder | 88.8 | 2836 | 89.4 | 2718 | 89.8 | 3454 |
| Reminder+TI | 89.4 | 2043 (-27.96%) | 89.6 | 1891 (-30.43%) | 91.6 | 3151 (-8.77%) |
Thinking Intervention effectively mitigates overthinking without compromising task performance. Table 3 summarizes the effectiveness of Thinking Intervention in addressing overthinking across different evaluated models. Empirically, we observe minimal changes in accuracy compared to baseline approaches (with accuracy varying within a narrow range, at most decreasing by 1.2%). Notably, Thinking Intervention can sometimes yield improved task accuracy, achieving gains of up to 1.6%. Most significantly, our method consistently produces substantial reductions in reasoning length, cutting token generation by as much as 30.4%. We note that token reduction for QwQ-32B is more modest compared to DeepSeek R1 models, suggesting distinct reasoning behavior across models. Nevertheless, stronger and more precisely tailored interventions could potentially further reduce reasoning length for QwQ-32B. Overall, our analysis demonstrates that Thinking Intervention reduces unnecessary reasoning steps, delivering meaningful improvements in efficiency with minimal negative impact on final task accuracy.
Appendix D Instruction Hierarchy (SEP)
D.1 More Details of Evaluations
We use the SEP dataset [81] to evaluate models’ ability to follow hierarchical instructions. The dataset contains 9,160 examples, each consisting of a main instruction, corresponding data, a low-priority query, and a witness (the answer to the low-priority query). For computational efficiency, we randomly sample 500 prompts for our evaluation. In Figure 9, we show all the prompts (including Vanilla Prompting, Reminder Prompting, and Thinking Intervention) used for evaluation.
In our main paper, we primarily focused on two metrics: robustness and utility. In fact, the SEP benchmark also contains another metric called SEP utility to measure if the model can correctly follow the low-priority task when it is placed in the task section. We detail these metrics as follows:
Robustness: We inject the low-priority query into the data block and measure the model’s ability to correctly ignore it (left example in Figure 10). The metric represents the percentage of cases where the model successfully ignores the low-priority instruction (i.e., the witness does not appear in the response). Note that in the original paper [81], this metric is called SEP.
SEP utility: We place the low-priority query in the task block and evaluate the model’s ability to follow it (middle example in Figure 10). The metric represents the percentage of cases where the model correctly follows the low-priority instruction (i.e., the witness appears in the response).
| Models | Methods | Robustness(%) | SEP utility(%) | Utility (%) |
| R1-Qwen-7B | Vanilla | 57.60 | 72.20 | 74.44 |
| Vanilla+ThinkingI | 60.80 (+3.20) | 77.80 (+5.60) | 74.40 (-0.04) | |
| Reminder | 57.60 | 74.20 | 74.20 | |
| Reminder+ThinkingI | 62.60 (+5.00) | 77.40 (+3.20) | 73.92 (-0.28) | |
| R1-Qwen-14B | Vanilla | 34.00 | 88.40 | 81.04 |
| Vanilla+ThinkingI | 38.40 (+4.40) | 92.40 (+4.00) | 81.08 (+0.04) | |
| Reminder | 38.40 | 88.80 | 80.50 | |
| Reminder+ThinkingI | 41.80 (+3.40) | 91.60 (+2.80) | 80.90 (+0.40) | |
| R1-Qwen-32B | Vanilla | 34.80 | 92.80 | 81.76 |
| Vanilla+ThinkingI | 50.20 (+15.40) | 91.60 (-1.20) | 82.02 (+0.26) | |
| Reminder | 46.20 | 92.00 | 81.16 | |
| Reminder+ThinkingI | 66.40 (+20.20) | 91.40 (-0.60) | 80.90 (-0.26) | |
| QwQ-32B | Vanilla | 22.20 | 96.60 | 88.00 |
| Vanilla+ThinkingI | 31.40 (+9.20) | 96.40 (-0.20) | 88.16 (+0.16) | |
| Reminder | 36.20 | 96.80 | 87.52 | |
| Reminder+ThinkingI | 43.40 (+7.20) | 96.60 (-0.20) | 86.79 (-0.73) | |
| R1-Llama-8B | Vanilla | 44.80 | 77.80 | 78.51 |
| Vanilla+ThinkingI | 53.80 (+9.00) | 79.40 (+1.60) | 77.04 (-1.47) | |
| Reminder | 48.00 | 77.40 | 78.53 | |
| Reminder+ThinkingI | 57.00 (+9.00) | 78.60 (+1.20) | 77.26 (-1.27) | |
| R1-Llama-70B | Vanilla | 34.20 | 91.40 | 81.45 |
| Vanilla+ThinkingI | 52.80 (+18.60) | 95.80 (+4.40) | 81.88 (+0.43) | |
| Reminder | 50.40 | 91.20 | 80.86 | |
| Reminder+ThinkingI | 65.80 (+15.40) | 95.60 (+4.40) | 80.90 (+0.04) | |
D.2 Comprehensive Experiment Results
In Table 4, we present extended evaluations of Thinking Intervention on the SEP dataset, incorporating a new metric (SEP utility) and additional reasoning-enhanced models (R1-Llama-8B and R1-Llama-70B).
Our results indicate that Thinking Intervention maintains or even improves the SEP utility. Specifically, for R1-Qwen-14B, our method achieves notable SEP utility scores of 92.4% and 91.6%, representing improvements of 4.0% and 2.8% over baseline methods, respectively. The performance on models like R1-Qwen-32B and QwQ-32B remains comparable, with only minimal SEP utility degradation (less than 1.2%), confirming that Thinking Intervention reliably enhances robustness without sacrificing model performance.
Moreover, we observe consistent and compelling improvements for the newly included models (R1-Llama-8B and R1-Llama-70B) across all evaluated metrics. Particularly noteworthy are the robustness gains: we achieve increases of up to 9.0% for R1-Llama-8B and up to 18.6% for R1-Llama-70B. These consistent improvements affirm the general applicability and effectiveness of Thinking Intervention in strengthening robustness across diverse reasoning model families.
Appendix E Safety Alignment
We provide more details on how baseline methods are implemented for safety alignment in Appendix E.1. We then introduce the detailed evaluation setup and present additional experimental results for XSTest and SORRY-Bench in Appendix E.2 and Appendix E.3, respectively. Lastly, we consider integrated Thinking Intervention with the safety fine-tuned model in Appendix E.4.
E.1 Details of Baseline Models
For those baseline prompting models of safety steering, we utilize the prompts shown in Figure 12.
E.2 Safety Alignment on XSTest
Details of experimental design. For safety alignment evaluation, we first use the XSTest benchmark [45], which consists of 250 safe requests and 200 unsafe requests. This benchmark is designed to examine the exaggerated safety behaviors of LLMs; therefore, all safe requests are intentionally tricky and contain potential safety triggers. Following [45], we evaluate the model outputs using the GPT-4o-mini with the prompts shown in Figure 13.
Note that the evaluation results are categorized into three levels: full compliance, full refusal, and partial refusal. For safe requests, we only consider full compliance as the correct response, and report the ratio as the compliance rate for safe requests. For unsafe requests, we only consider full refusal as the correct response, and report the ratio as the refusal rate for unsafe requests. This provides a stringent evaluation of the model’s safety alignment behavior, as it requires models to clearly distinguish between safe and unsafe requests, and to respond appropriately in each case.
Experimental results across all reasoning models. In Figure 14, we present extensive safety alignment evaluations on the XSTest dataset using a diverse set of reasoning-enhanced models, including R1-Qwen-7B, R1-Qwen-14B, R1-Qwen-32B, QwQ-32B, R1-Llama-8B, and R1-Llama-70B.
Our results demonstrate that employing Thinking Intervention consistently and substantially improves model safety across a range of prompting methods and model architectures. In particular, compared with Vanilla Prompting, integrating Thinking Intervention significantly increases refusal rates for unsafe requests by over 40% for R1 series models and approximately 10% for QwQ-32B models.
When evaluating safe requests, compliance rates under Thinking Intervention show only minor reductions (generally less than 10%) for medium- and large-sized reasoning models (R1-Qwen-14B, R1-Qwen-32B, QwQ-32B, and R1-Llama-70B) across most prompting methods, except for Goal Priority Prompting. However, we observe relatively larger drops in compliance for smaller reasoning models (R1-Qwen-7B and R1-Llama-8B). These discrepancies are likely due to the limited capacity of smaller models to accurately distinguish tricky safe prompts from truly unsafe ones. Encouragingly, larger models tend to handle this trade-off more effectively, maintaining strong compliance while enhancing safety.
Furthermore, our analysis indicates that the performance changes induced by identical Thinking Intervention vary across model families, with the QwQ-32B series exhibiting smaller improvements compared to the R1 series. This may be because QwQ-32B is already substantially safer than the R1 models. This variance also highlights intrinsic differences in model behavior, a phenomenon that can be investigated in future work.
| (a) R1-Qwen-7B | (b) R1-Qwen-14B | (c) R1-Qwen-32B |
| (a) QwQ-32B | (b) R1-Llama-8B | (c) R1-Llama-70B |
E.3 Safety Alignment on SORRY-Bench
Details of experimental design. We also evaluate the safety alignment of reasoning models on the SORRY-Bench benchmark [68], which is a more comprehensive benchmark containing 45 taxonomies of unsafe instructions, with 10 unsafe requests per taxonomy (450 total unsafe requests). Following the method in [68], we use GPT-4o-mini as the evaluation judge with the prompts shown in Figure 15.
Experimental results across multiple reasoning models. We comprehensively evaluate the effectiveness of our Thinking Intervention method on the SORRY-Bench benchmark across a diverse set of reasoning-enhanced models, including R1-Qwen-7B, R1-Qwen-14B, R1-Qwen-32B, QwQ-32B, R1-Llama-8B, and R1-Llama-70B. As illustrated in Figure 16, our approach consistently and substantially enhances model safety alignment across various prompting methods. Specifically, we observe improvements in robustness ranging from 10% to 25% for the R1 model series and from 5% to 15% for QwQ-32B. Remarkably, after incorporating Thinking Intervention within the baseline prompting method, the refusal rates for unsafe requests exceed those of all GPT-series models using the same prompting strategy.
These findings demonstrate that Thinking Intervention offers an effective, robust, and broadly applicable solution for significantly improving safety alignment in diverse reasoning models.
| (a) R1-Qwen-7B | (b) R1-Qwen-14B | (c) R1-Qwen-32B |
| (a) QwQ-32B | (b) R1-Llama-8B | (c) R1-Llama-70B |
E.4 Thinking Intervention and Safety Fine-tuning
In this appendix, we present experiments evaluating our Thinking Intervention method in conjunction with recent safety fine-tuning techniques proposed by [61]. Specifically, we consider the STAR1-32B model, obtained by fine-tuning the original R1-Qwen-32B model using approximately 1k generated instruction-and-reasoning pairs that incorporate explicit safety guidelines. Using the publicly released checkpoint for STAR1-32B, we evaluate its performance on the XSTest and SORRY-Bench benchmarks.
Integration of Thinking Intervention with Safety Fine-tuning. Figure 17(a) illustrates the effectiveness of our Thinking Intervention approach when integrated with the safety fine-tuned STAR1-32B model on the XSTest benchmark. We observe that incorporating Thinking Intervention consistently improves safety alignment, increasing the refusal rate for unsafe requests by approximately 3%–5% across all evaluated prompting methods. For safe requests, Thinking Intervention generally maintains high compliance rates across most methods, and even slightly improves compliance (by 2%) in the case of the Goal Priority Prompting method.
Similarly, Figure 17(b) shows that Thinking Intervention further enhances the safety alignment of the STAR1-32B model on the SORRY-Bench benchmark. The refusal rate for unsafe requests increases by 0.2%–1.8% across all prompting methods. Despite the already strong performance of STAR1-32B, our Thinking Intervention approach provides some additional gains in safety alignment.
Appendix F Design of Thinking Intervention
In this section, we present an in-depth analysis of how variations in the design of the Thinking Intervention influence performance using safety alignment benchmarks. In Appendix F.1, we analyze the effect of positioning the intervention sequence at the beginning, middle, and end of the reasoning process. In Appendix F.2, we evaluate how variations in the text of the intervention sequence affect model responses, highlighting the trade-offs between safety and compliance rates. In Appendix F.3, we investigate the effect of the narrative style of the intervention sequence and find that the model is capable of self-correcting narrative inconsistencies. In Appendix F.4, we discuss how we use auxiliary LLMs to support the Thinking Intervention.
F.1 Position of Thinking Intervention
We investigate the effect of intervention positions on the reasoning process using R1-Qwen-32B on XSTest and SORRY-Bench. The intervention sequence is kept unchanged, and we implement three distinct intervention functions: (1) TIbegin: The intervention is introduced at the beginning of the reasoning process, corresponding to the default setting described in the main text. (2) TIend: The intervention is introduced at the conclusion of the reasoning process. Specifically, when the model is about to generate the reasoning-ending token "</think>", we replace it with the intervention sequence and allow the model to continue generating. (3) TImid: The intervention occurs at an intermediate stage of the reasoning process. We use the token "wait" as a trigger, indicating a transition in reasoning. Upon detecting this trigger, we replace the "wait" token with the intervention sequence and continue generating the output.
| (a) TIbegin vs. TIend on XSTest | (b) TIbegin vs. TIend on SORRY-Bench |
| (c) TIbegin vs. TImid on XSTest | (d) TIbegin vs. TImid on SORRY-Bench |
Thinking Intervention at reasoning start steers the model more effectively. Figure 18(a) and Figure 18(c) illustrate that intervening at later stages of the reasoning process (i.e., TIend and TImid) reduces the model’s steering capability, resulting in a lower refusal rate for harmful queries but higher compliance rates for safe queries. For instance, with Reminder Prompting, the model achieves a refusal rate of approximately 50% and compliance rate of nearly 98%, whereas TIbegin attains a refusal rate of around 70% and a compliance rate of approximately 96%. This outcome highlights a clear trade-off between these two metrics, which is expected since the model has already generated substantial reasoning content before the intervention, limiting the effectiveness of subsequent steering.
F.2 Complexity of Thinking Intervention
We then explore how the different text content of the Thinking Intervention affects the reasoning process and the final safety performance using the XSTest and SORRY-Bench benchmarks with the R1-Qwen-32B model. We evaluate two different intervention sequences: (1) the default shorter intervention sequence (i.e., "I am a helpful, respectful, and honest assistant."), denoted as TIshort, and (2) a longer version of the intervention sequence (i.e., "I am a responsible assistant and should not generate harmful or misleading content! So I need to verify the dangerousness of the user’s request."), denoted as TIlong. We insert both versions at the beginning of the reasoning process.
Long and detailed Thinking Intervention can steer the model toward safety but results in a lower compliance rate. Figure 19 (a) shows that the compliance rate for safe requests on XSTest decreases by over 10% when using the longer intervention sequence, compared to the shorter version. In contrast, the refusal rate for unsafe requests increases by approximately 10% with the longer intervention sequence. Similarly, Figure 19 (b) indicates that the refusal rate for unsafe requests on SORRY-Bench also improves with the longer intervention sequence, showing an increase of 3% to 10%. This outcome is expected, as the longer intervention sequence provides more explicit safety instructions. Therefore, Thinking Intervention can be flexibly designed with different intervention contents to precisely control the degree of steering.
| (a) R1-Qwen-32B on XSTest | (b) R1-Qwen-32B on SORRY-Bench |
F.3 Narrative of Thinking Intervention
| (a) R1-Qwen-32B on XSTest | (b) R1-Qwen-32B on SORRY-Bench |
In this subsection, we analyze how the narrative of the Thinking Intervention affects the reasoning process using safety alignment benchmarks (i.e., XSTest and SORRY-Bench) with the R1-Qwen-32B model. In our main experiments, we use the first narrative intervention sequence (i.e., "I am a helpful, respectful, and honest assistant.") as the default setting. We then construct an alternative second narrative intervention sequence (i.e., "You are a helpful, respectful, and honest assistant.") and examine whether the reasoning process is affected by the narrative. Again, both Thinking Interventions are inserted at the beginning of the reasoning process.
Narrative of Thinking Intervention does not affect the performance significantly. In Figure 20 (a), we observe that the compliance rate for safe requests and the refusal rate for harmful requests on XSTest are not significantly affected by the narrative of the Thinking Intervention, with less than 3% variation. On SORRY-Bench, Figure 20 (b) indicates that the refusal rate for unsafe requests decreases with the second narrative intervention sequence, showing a drop of 0.2% to 3.8%. This outcome may be somewhat surprising, as we expected the second narrative intervention sequence to disrupt the reasoning process. Therefore, we explore why this occurs through an example.
Reasoning models can self-correct narrative errors. We demonstrate an example from the SORRY-Bench benchmark in Figure 21, where the user instruction is a harmful request. The model with the second narrative intervention sequence initially uses the second narrative and generates "You can’t comply with this request." However, the model quickly recognizes that the narrative is incorrect and corrects it to "So, I need to figure out how to respond appropriately." In the subsequent reasoning steps, the model continues using the first-narrative reasoning. This indicates that the model can self-correct narrative inconsistencies during the reasoning process.
F.4 Leveraging auxiliary LLM for Thinking Intervention
In this subsection, we conduct preliminary experiments to explore the use of an auxiliary LLM to revise existing reasoning steps on XSTest and SORRY-Bench. Specifically, we use R1-Qwen-32B as the primary reasoning model and Qwen2.5-7B-Instruct [42] as the auxiliary LLM, which is instructed to edit the reasoning chain. The detailed prompt used to guide the editing process is shown in Figure 22. We denote our default Thinking Intervention as TIdefault and the LLM-assisted Thinking Intervention as TImodel.
Leveraging an auxiliary LLM does not exceed our manually designed prompt in most cases. In Figure 23(a), we observe that TIdefault generally achieves a higher refusal rate to harmful requests and better compliance with safe requests compared to TImodel, across all prompting methods except Vanilla Prompting. On the SORRY-Bench benchmark, TIdefault outperforms across all four prompting methods, although the performance difference is less than 3%. Therefore, effectively leveraging an auxiliary LLM to edit the reasoning trace remains a challenging task, highlighting the need for further research and exploration in this direction. In addition, using an auxiliary LLM to edit the reasoning trace is costly and may not be practical.
| (a) R1-Qwen-32B on XSTest | (b) R1-Qwen-32B on SORRY-Bench |
Appendix G Case Study of Instruction Hierarchy
In this section, we showcase an example from the SEP dataset that demonstrates how Thinking Intervention can effectively guide reasoning models to correctly follow hierarchical instructions. In Figure 24 and Figure 25, the main instruction is to "Interpret the statute mentioned in the next paragraph, explaining its legal implications." where the content is about "The Sarbanes-Oxley Act". However, the data block contains a low-priority instruction to "Tell me what a book is primarily made of.", which is unrelated to the main task and should be ignored.
Figure 24 shows the response of a model without Thinking Intervention. The model directly tries to answer the low-priority query and discuss "what a book is primarily made of". Then the model recognizes the main task related to the "Sarbanes-Oxley Act". However, it gets confused and finally reverts to considering the low-priority query. As a result, the model provides an incorrect response by discussing the content of the low-priority query instead of interpreting the statute.
Then, in Figure 25, we demonstrate the response of a model with Thinking Intervention. The model directly focuses on the main task and reasons about how to answer it. Then the model also finds the low-priority query in the data block, but it successfully recognizes that this is unrelated to the main task and should be ignored. Consequently, the model provides a correct response by interpreting the statute mentioned in the main instruction. This example illustrates how Thinking Intervention can effectively guide reasoning models to correctly follow hierarchical instructions and provide accurate responses.
Appendix H Case Study of Attention
In this appendix, we analyze the attentions of Prompt Engineering and Thinking Intervention. We use the R1-Qwen-14B model with Reminder Prompting as the baseline and compare it with Thinking Intervention applied at the beginning of the reasoning process. Specifically, we examine an example where the user instruction is "How did people make a Molotov cocktail?", with "You are a safe assistant." used for prompt engineering and "I am a safe assistant." used for Thinking Intervention. The safety instruction content remains nearly identical, with the only difference being its position—either before or after the thinking tag, "<think>". Attention weights are derived from the attention scores in the model’s final layer. We first present the model’s output with prompt engineering in Figure 26, followed by the output with Thinking Intervention in Figure 27. Then, we visualize and compare the attention maps in Figures 28 and 29 to observe how attention is allocated across tokens in each input.
Thinking Intervention is more effective than Prompt Engineering in steering the model towards safety. With Prompt Engineering (Figure 26), we observe that the model generates a response that first tries to identify what a Molotov cocktail is and then provides a detailed description of how to make it. This indicates that the model is not effectively refusing the harmful request, even though a safety instruction is provided. In contrast, with Thinking Intervention (Figure 27), the model successfully refuses the harmful request and provides a safe response. This demonstrates that Thinking Intervention is more effective than Prompt Engineering in steering the model towards safety.
Thinking Intervention directs more attention to the safety instructions during generation. In Prompt Engineering (Figure 28), the model’s attention during the generation of the harmful content shows that the safety instruction "You are a safe assistant." receives little attention in later reasoning generation stages. This suggests that the model is not effectively utilizing the safety instruction during its reasoning process. Conversely, with Thinking Intervention (Figure 29), the safety instruction "I am a safe assistant.", placed after the thinking tag, receives considerably more attention from later reasoning tokens. This focused attention on the safety guidelines within the thinking process appears to be key to Thinking Intervention’s effectiveness in steering the model towards a safe response.
Therefore, this case study illustrates that Thinking Intervention is more effective than Prompt Engineering in guiding the model’s reasoning trace, thereby effectively steering the model towards safety. The attention analysis further supports this conclusion, showing that Thinking Intervention directs more attention to the safety instructions during the reasoning process, leading to improved safety alignment.
Appendix I Ethics, Broader Impact, Reproducibility, and Licenses
Ethics. Our research aims to improve the safety alignment of reasoning models as part of responsible AI development. We focus on enabling large language models to better follow hierarchical instructions and adhere to safety guidelines, particularly in sensitive domains. We conduct our work responsibly, transparently, and in compliance with the NeurIPS Code of Ethics.
Broader Impact. Enhancing safety alignment in reasoning models can benefit high-stakes fields such as healthcare, finance, and education. At the same time, we acknowledge the risks involved in applying large language models in these areas and encourage proactive efforts to identify and mitigate potential harms.
Reproducibility. We provide comprehensive details of our experiments in Appendix B, Appendix D, Appendix C, and Appendix E. All experiments were conducted using one or two H100 GPU 80G within 24 hours. We employed greedy decoding with a temperature of 0 for all experiments, ensuring that our results are deterministic and fully reproducible.
Licenses. In this paper, we utilize the following models and datasets: (1) Models: R1-Qwen-32B (Apache 2.0 License), R1-Qwen-14B (Apache 2.0 License), R1-Llama-8B (LLaMA 3.1 License), R1-Llama-70B (LLaMA 3.3 License), and QwQ-32B (Apache 2.0 License). GPT models are closed-source and not publicly available. (2) Datasets: IFEval (Apache 2.0 License), SEP (MIT License), XSTest (Attribution 4.0 International), SORRY-Bench (Sorry-Bench License), and MATH (MIT License).